OpenStudy (anonymous):

Find the Discriminant. x^2-30x-64=0

OpenStudy (perl):

The discriminant is defined as $$\Large{ ax^2 +bx + c = 0 \\~\\ \Delta = b^2 - 4ac }$$

OpenStudy (kidrah69):

\[y=2x^2−30x−64 \] \[discriminant=b^2−4(a)(c)\] \[(−30)2−4(2)(−64)=? \]

OpenStudy (er.mohd.amir):

ax2+bx+c=0 diecriment D=sqr(b2-4ac) put value and solve

OpenStudy (anonymous):

-132?

OpenStudy (perl):

not quite

OpenStudy (kidrah69):

(−30)^2−4(2)(−64)

OpenStudy (kidrah69):

oops sorry :/ must've made a typo... yeah ignore the 2 @yoda5657

OpenStudy (kidrah69):

(-30)^2 - 4(1)(-6 <---solve

OpenStudy (anonymous):

I'm getting 1156.

OpenStudy (kidrah69):

thats correct :)

OpenStudy (anonymous):

Thanks :)

OpenStudy (kidrah69):

thank @perl too :)

OpenStudy (perl):

:)